A solution of 35% glycol is held in a 400 gallon tank. A 72% solution of glycol is held in a 190 gallon tank. How much of the two tanks must be added together to have a 200 gallon tank of 50% glycol?
Answer by nerdybill(7384)

A solution of 35% glycol is held in a 400 gallon tank. A 72% solution of glycol is held in a 190 gallon tank. How much of the two tanks must be added together to have a 200 gallon tank of 50% glycol?
.
Let x = amount (gallons) of 35%
then
200-x = amount (gallons) of 72%
.
.35x + .72(200-x) = .50(200)
.35x + 144-.72x = 100
144-.37x = 100
-.37x = -44
x = 118.9 gallons (35% solution)
.
72% solution:
200-x = 200-118.9 = 81.1 gallons